Thanks for your considration. A six passenger plane cruises at 180mph in calm air. if the plane flies 7miles with the wind in the same amount of time as it flies 5miles against the wind, then what is the wind speed?
Found 2 solutions by stanbon, edjones:
Answer by stanbon(75887)

A six passenger plane cruises at 180 mph in calm air. if the plane flies 7 miles with the wind in the same amount of time as it flies 5 miles against the wind, then what is the wind speed?
-------------
Downwind DATA:
distance = 7 mi ; rate = 180+w ; time = d/t = 7/(180+w)
---------------------
Upwind DATA:
distance = 5 mi ; rate = 180-w ; time = d/t = 5(180-w)
---------------------
Equation:
time = time
7/(180+w) = 5/(180-w)
7*180 - 7w = 5*180 + 5w
12w = 2*180
w = 30 mph (speed of the wind)
---------------------------------
Cheers,
Stan H.



Answer by edjones(8007)

d=st
t=d/s
s=180
Let the wind speed be w
because the time it takes is equal d/s =d/s
7/(180+w) = 5/(180-w)
7(180-w)=5(180+w)
1260-7w=900+5w
12w=360
w=30mph
.
Ed